OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely.

OpenStudy (anonymous):

\[7\sqrt{a}\]

OpenStudy (anonymous):

Do you always talk to yourself?

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

uhh what?

OpenStudy (anonymous):

is there an easier explanation as what the heck your learning?

OpenStudy (anonymous):

Simplifying Radical Expressions

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

its friggin algebra 1 lol

OpenStudy (anonymous):

ahhhhhhhh i FAILED algebra 1

OpenStudy (anonymous):

lol jk jk bro I'm actually in it right now

OpenStudy (anonymous):

u do this yet?

OpenStudy (anonymous):

I only passed bc they let me retake a test in my first semester. I finished with a D- the first sem and D+ the second

OpenStudy (anonymous):

WOW i finished with a C on my first and now i think i might finish with a C- on my last not much difference lol

OpenStudy (anonymous):

I'm in Alg II with a C+

OpenStudy (anonymous):

i have an 80% cause of OS lol

OpenStudy (anonymous):

wow brad you little cheapskape

OpenStudy (anonymous):

whatevs bro

OpenStudy (anonymous):

lol xD

OpenStudy (anonymous):

I suggest you find somebody who DOESN'T suck at algebra to help you cuz if i help you this will end up looking like a polynomial

OpenStudy (anonymous):

dang it mud who here is smart?

OpenStudy (anonymous):

found answer dont bother with it any1 :)

OpenStudy (anonymous):

\[\sqrt{125x} = 5\sqrt{5x}\]

OpenStudy (anonymous):

if you want one that simplifies to \[7\sqrt{a}\] then \[\sqrt{49a} = 7\sqrt{a}\]